# The effects of star-gas interactions on binary evolution in open clusters

### v1

@@ IV Binary dynamics in clusters immersed in gas
To ensure computational efficiency and avoid the omission of wide pairs, we identify candidate binaries with a maximum separation $s_{\mathrm{max}}=5~\mathrm{pc}$ (Section II.4), followed by our standard boundness criteria. As the separation of widest binary in the cluster is $\sim 10^{4}~{\rm AU}\sim 0.05~{\rm pc}$ (2003gmbp.book.....H; 2020ApJ...904..113R), we have also tested $s_{\mathrm{max}}=0.1~\mathrm{pc}$ and obtained very similar results, indicating that our conclusions are not sensitive to this choice.
@@
We classify binaries by their binding energy, $E_{\rm b}=-E_{\rm tot}>0$, where $E_{\rm tot}$ is defined in Eq. (11). Hard binaries are defined as systems whose binding energy exceeds the local average kinetic energy of stars in the cluster ($E_{\rm b}>\overline{E}_{\rm k}$), whereas soft binaries have binding energies below the local average kinetic energy ($0<E_{\rm b}<\overline{E}_{\rm k}$) and are therefore easily disrupted by encounters. Heggie’s law (1975MNRAS.173..729H) states that hard binaries tend to become harder and soft binaries tend to become softer as a result of multi-body encounters. DF from the ambient gas is expected to increases the stellar density and encounter frequency, which likely enhances this effect. In contrast, NDF leads to a more diffuse stellar distribution, which may reduce the encounter rate and is consistent with an increased survival probability of soft binaries.
@@
We also analyze the orbital parameters of the surviving binaries, including eccentricity $e$, semi-major axis $a$, total mass $M_{\rm tot}$ and orbital inclination $i$. The distribution of $\cos{i}$ is essentially uniform, which indicates that the cluster remains nearly spherically symmetric throughout its evolution. e,ae,~a and $M_{\rm tot}$ are shown in Figure 13. Results of NDF-D3E6 and NDF-D3E8 show that NDF tends to preserve binaries with larger semi-major axes, producing wider systems, decreasing the binding energy of binary systems. This is demonstrated in the study of a separate binary system (2022ApJ...932..108W) that outflows in binary systems exert positive torques and expand the orbits. Binaries in NDF-D3E8 also exhibit smaller $e$ and smaller $M_{\rm tot}$ than those in other models. This reflects a preferential disruption of binaries containing more massive stars, which experience stronger non-dissipative interactions in our model. As a result, the surviving binary population is biased toward more circular orbits and lower total masses, reducing the likelihood of stellar collisions and mergers.
@@ V Discussion and summary
The interstellar medium densities explored span from typical Milky Way conditions ($30~m_{p}~\mathrm{cm^{-3}}$) to extreme environments such as AGN disks ($3\times 10^{8}~m_{p}~\mathrm{cm^{-3}}$). While the latter may not represent typical OC habitats, these high-density simulations serve as controlled scaling experiments that clarify the dependence of gas-star interaction effects on ambient density. The low- to intermediate-density regimes (3030-$300~m_{p}~\mathrm{cm^{-3}}$), which are more relevant for typical OCs, already demonstrate that gas interactions can significantly influence binary evolution timescales and population demographics.
Our results are broadly consistent with cluster-scale dynamical evolution reported by 2025ApJ...978...87L. An important improvement of the present study is the more robust numerical treatment of the binaries. The relatively low number of surviving binaries in 2025ApJ...978...87L likely reflects the combined effects of smaller particle numbers, which amplify stochasticity, and the use of integration schemes not optimized for long-term resolution of close encounters. In contrast, PeTar is specifically designed for cluster-scale simulations, with accurate handling of binaries, triples, and higher-order systems, leading to improved overall numerical reliability. Consistent with 2025ApJ...978...87L, we find no evidence that NDF universally suppresses the binary population, nor that DF leads to a substantial binary enhancement over the simulation timescales. Our simulations extend these conclusions by resolving how gas-star interactions reshape binary orbital parameters and subtype demographics.
@@
Finally, our treatment assumes isotropic outflows, whereas real stellar winds and jets are often anisotropic. Li_NDF_2020 showed that strongly collimated outflows can suppress or eliminate the NDF effect by preventing the formation of a symmetric underdense cavity. In such cases, the net force remains decelerating, though reduced relative to the standard DF case. The NDF effects reported here should therefore be regarded as upper limits, and the impact of anisotropic outflows warrants dedicated future study.
Although direct observational constraints on the survival boundary of wide and dynamically soft binaries in open clusters remain limited, several lines of evidence suggest that wide systems can exist under specific environmental conditions. High-angular-resolution imaging surveys have revealed the presence of massive companions at separations up to $\sim 10^{3}$$10^{4}\,\mathrm{AU}$ in young associations such as Cyg OB2 (2014AJ....147...40C). More generally, multiplicity surveys of nearby star-forming regions show substantial populations of wide systems over separations of $\sim 3$$5000\,\mathrm{AU}$, providing an empirical baseline for the initial wide-binary inventory prior to significant dynamical processing (e.g., Kraus2011). These observations suggest that environmental effects beyond simple stellar-encounter rates may play a role in regulating the survival of soft binaries (2007ApJ...662..413K).
@@ Acknowledgments
The computation resource supporting this work is provided by the KIAA. We thank our colleagues Long Wang and Chuizheng Kong for helpful discussions. We are also grateful to Sanaea Rose for insightful discussions on the dynamical definitions and classification of binary systems.

### v2

@@ IV Binary dynamics in clusters immersed in gas
To ensure computational efficiency and avoid the omission of wide pairs, we identify candidate binaries with a maximum separation $s_{\mathrm{max}}=5~\mathrm{pc}$ (Section II.4), followed by our standard boundness criteria. As the separation of widest binary in the cluster is $\sim 10^{4}~{\rm AU}\sim 0.05~{\rm pc}$ (Heggie and Hut, 2003; Rose et al., 2020), we have also tested $s_{\mathrm{max}}=0.1~\mathrm{pc}$ and obtained very similar results, indicating that our conclusions are not sensitive to this choice.
@@
We classify binaries by their binding energy, $E_{\rm b}=-E_{\rm tot}>0$, where $E_{\rm tot}$ is defined in Eq. (11). Hard binaries are defined as systems whose binding energy exceeds the local average kinetic energy of stars in the cluster ($E_{\rm b}>\overline{E}_{\rm k}$), whereas soft binaries have binding energies below the local average kinetic energy ($0<E_{\rm b}<\overline{E}_{\rm k}$) and are therefore easily disrupted by encounters. Heggie’s law (Heggie, 1975) states that hard binaries tend to become harder and soft binaries tend to become softer as a result of multi-body encounters. DF from the ambient gas is expected to increases the stellar density and encounter frequency, which likely enhances this effect. In contrast, NDF leads to a more diffuse stellar distribution, which may reduce the encounter rate and is consistent with an increased survival probability of soft binaries.
@@
We also analyze the orbital parameters of the surviving binaries, including eccentricity $e$, semi-major axis $a$, total mass $M_{\rm tot}$ and orbital inclination $i$. The distribution of $\cos{i}$ is essentially uniform, which indicates that the cluster remains nearly spherically symmetric throughout its evolution. e,ae,~a and $M_{\rm tot}$ are shown in Figure 13. Results of NDF-D3E6 and NDF-D3E8 show that NDF tends to preserve binaries with larger semi-major axes, producing wider systems, decreasing the binding energy of binary systems. This is demonstrated in the study of a separate binary system (Wang and Li, 2022) that outflows in binary systems exert positive torques and expand the orbits. Binaries in NDF-D3E8 also exhibit smaller $e$ and smaller $M_{\rm tot}$ than those in other models. This reflects a preferential disruption of binaries containing more massive stars, which experience stronger non-dissipative interactions in our model. As a result, the surviving binary population is biased toward more circular orbits and lower total masses, reducing the likelihood of stellar collisions and mergers.
@@ V Discussion and summary
The interstellar medium densities explored span from typical Milky Way conditions ($30~m_{p}~\mathrm{cm^{-3}}$) to extreme environments such as AGN disks ($3\times 10^{8}~m_{p}~\mathrm{cm^{-3}}$). While the latter may not represent typical OC habitats, these high-density simulations serve as controlled scaling experiments that clarify the dependence of gas-star interaction effects on ambient density. The low- to intermediate-density regimes ($30$-$300~m_{p}~\mathrm{cm^{-3}}$), which are more relevant for typical OCs, already demonstrate that gas interactions can significantly influence binary evolution timescales and population demographics.
Our results are broadly consistent with cluster-scale dynamical evolution reported by Liu et al. (2025). An important improvement of the present study is the more robust numerical treatment of the binaries. The relatively low number of surviving binaries in Liu et al. (2025) likely reflects the combined effects of smaller particle numbers, which amplify stochasticity, and the use of integration schemes not optimized for long-term resolution of close encounters. In contrast, PeTar is specifically designed for cluster-scale simulations, with accurate handling of binaries, triples, and higher-order systems, leading to improved overall numerical reliability. Consistent with Liu et al. (2025), we find no evidence that NDF universally suppresses the binary population, nor that DF leads to a substantial binary enhancement over the simulation timescales. Our simulations extend these conclusions by resolving how gas-star interactions reshape binary orbital parameters and subtype demographics.
@@
Finally, our treatment assumes isotropic outflows, whereas real stellar winds and jets are often anisotropic. Li et al. (2020) showed that strongly collimated outflows can suppress or eliminate the NDF effect by preventing the formation of a symmetric underdense cavity. In such cases, the net force remains decelerating, though reduced relative to the standard DF case. The NDF effects reported here should therefore be regarded as upper limits, and the impact of anisotropic outflows warrants dedicated future study.
Although direct observational constraints on the survival boundary of wide and dynamically soft binaries in open clusters remain limited, several lines of evidence suggest that wide systems can exist under specific environmental conditions. High-angular-resolution imaging surveys have revealed the presence of massive companions at separations up to $\sim 10^{3}$$10^{4}\,\mathrm{AU}$ in young associations such as Cyg OB2 (Caballero-Nieves et al., 2014). More generally, multiplicity surveys of nearby star-forming regions show substantial populations of wide systems over separations of $\sim 3$$5000\,\mathrm{AU}$, providing an empirical baseline for the initial wide-binary inventory prior to significant dynamical processing (e.g., Kraus et al., 2011). These observations suggest that environmental effects beyond simple stellar-encounter rates may play a role in regulating the survival of soft binaries (Kraus and Hillenbrand, 2007).
@@ Acknowledgments
The computational resources supporting this work are provided by the Kavli Institute for Astronomy and Astrophysics (KIAA).
Special thanks are extended to Chao Lin for his contributions to the development of the dynamical friction module in PeTar. We thank our colleagues Huan Yang and Chuizheng Kong for helpful discussions, and we are also grateful to Sanaea Rose for insightful discussions on the dynamical definitions and classification of binary systems.
L.W. acknowledges support from the National Natural Science Foundation of China (Grant Nos. 12233013 and 12573041), the High-level Youth Talent Project (Provincial Financial Allocation; Grant No. 2023HYSPT0706), and the Fundamental Research Funds for the Central Universities at Sun Yat-sen University (Grant No. 2025QNPY04).
M.S. acknowledges support from the National Natural Science Foundation of China (NSFC) through the Fundamental Science Center for Nearby Galaxies (Grant No. 12588202), a project based on LAMOST and FAST, as well as from an additional NSFC grant (Grant No. 12503041; PI: Meng Sun). M.S. also acknowledges support from the Gordon and Betty Moore Foundation through Grant GBMF8477 (PI: Vicky Kalogera).
R.S. acknowledges support from the German Research Foundation (DFG; Grant No. Sp 345/24-1) and from the NSFC (Grant No. 12473017). The International Cooperation Office of the National Astronomical Observatories of the Chinese Academy of Sciences (NAOC, Beijing) supported the work of R.S. in China during 2023–2025. Since 2026, support from the Chinese Academy of Sciences President’s International Fellowship Initiative for Visiting Scientists (PIFI; Grant No. 2026PVA0089) is acknowledged.